OpenStudy (gatsbyyy):

A certain recipe requires 2/9 cup of flour and 3/10 cup of sugar. A. If 5/8 of the recipe is to be made, how much sugar is needed? Write your answer as a proper or improper fraction. B. If the above ingredients are require for one batch, find the amount of flour needed for a triple batch. Write your answer as a proper or improper fraction.

OpenStudy (hannaha):

sorry i have no help, just wondering who in the heck cooks with 2/9 cup of flour and 3/10 cup of sugar XP for real

satellite73 (satellite73):

five eighths of three tenths means \[\frac{5}{8}\times \frac{3}{10}\]

satellite73 (satellite73):

cancel first multiply last